# VR-Doh: Hands-on 3D Modeling in Virtual Reality

### v2

@@ Abstract.
We introduce VR-Doh, a hands-on 3D modeling system that enables intuitive creation and manipulation of elastoplastic objects in Virtual Reality (VR). By customizing the Material Point Method (MPM) for real-time simulation of hand-induced large deformations and enhancing 3D Gaussian Splatting for seamless rendering, VR-Doh provides an interactive and immersive 3D modeling experience. Users can naturally sculpt, deform, and edit objects through both contact- and gesture-based hand-object interactions. To achieve real-time performance, our system incorporates localized simulation techniques, particle-level collision handling, and the decoupling of physical and appearance representations, ensuring smooth and responsive interactions. VR-Doh supports both object creation and editing, enabling diverse modeling tasks such as designing food items, characters, and interlocking structures, all resulting in simulation-ready assets. User studies with both novice and experienced participants highlights the system’s intuitive design, immersive feedback, and creative potential. Compared to existing geometric modeling tools, VR-Doh offers enhanced accessibility and natural interaction, making it a powerful tool for creative exploration in VR.
@@
submissionid: 318journalyear: 2025isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Virtual reality
@@ 1. Introduction
Meanwhile, 3D modeling in Virtual Reality (VR) has become increasingly popular due to its ability to provide immersion and depth, thereby enhancing designers’ creativity and collaboration (Rosales et al., 2019; Yu et al., 2021). Commercial tools like Shapelab (sha, 2025) and Gravity Sketch (gra, 2025) have advanced this vision. However, their approaches are still limited to procedural methods, such as drawing curves to form surfaces from scratch, which are restricted to geometric techniques and require significant skills.
@@
Building on these foundational advantages, we integrate physics-based simulation into VR to deliver an intuitive and efficient hands-on 3D modeling system, VR-Doh. Our primary objective is to enable users to create and edit virtual objects with a high degree of realism, leveraging direct hand-based contact and gestures alongside interactive feedback. To efficiently support large deformations and contact handling, we adopt the Material Point Method (MPM) (Jiang et al., 2016; Hu et al., 2018), a hybrid Lagrangian-Eulerian framework capable of simulating versatile elastoplastic materials.
VR-Doh enables users to create 3D models by deforming and composing built-in primitive geometries, with surface meshes reconstructed from MPM particles. Additionally, based on Xie et al. (2024), we extend support for editing existing 3D data represented by Gaussian Splatting (GS) (Kerbl et al., 2023), allowing content creation through the modification of realistically rendered objects captured from the real world. To achieve real-time responsiveness despite the computational demands of simulation and rendering, we introduce optimizations such as particle-level collision handling and the decoupling of appearance and physical representations, allowing detailed hand-object interactions within budgeted simulation degrees of freedom. Beyond hand-based contact and gestures, we integrate a suite of deformation tools driven by tracked hand motions, providing diverse and precise deformations to enhance users’ creative flexibility. To evaluate VR-Doh’s usability and effectiveness, we conducted user studies with participants of varying expertise. The study highlights its intuitive design, immersive feedback, and potential for complementing traditional modeling software in creative exploration. These findings offer insights into how users naturally leverage physics-based interactions, paving the way for future improvements and more immersive 3D modeling experiences in VR.
@@
A key advantage of our approach is its ability to make 3D modeling in VR significantly more intuitive and accessible, especially for novice users. By incorporating physics-based deformations, VR-Doh provides a realistic and immersive modeling experience that surpasses traditional geometric modeling approaches. Hands-on input lowers the barrier to entry, enabling users to interact with virtual objects naturally and intuitively, while flexible selection of operation areas improves efficiency and enhances immersion. This direct interaction mimics real-world manipulation, allowing users to predict outcomes based on everyday experiences and reducing the reliance on specialized skills and technical expertise.
@@ 2. Related Work
Virtual modeling of clay-like materials using dexterous input offers significant advantages over traditional 3D modeling, including high expressivity and a lower learning curve, effectively replicating the tactile experience of working with physical clay (Sheng et al., 2006; Chatterjee, 2024). Despite these benefits, the computational demands of simulating dexterous hand-object interactions have limited the integration of virtual clay modeling into VR environments, with most existing implementations constrained to 2D screens. Early approaches to virtual clay modeling utilized dynamic subdivision solids (McDonnell et al., 2001), followed by the application of plasticity models to capture essential physical properties of real clay, such as mass conservation and surface tension effects (Dewaele and Cani, 2004). To improve interactivity, Barreiro et al. (2021) introduced a particle-based viscoplasticity model with ultrasound haptic feedback, although the fidelity of finger-based interactions remained insufficient for detailed 3D modeling tasks. Additionally, physical proxies have been employed to enhance the virtual clay experience (Sheng et al., 2006; Marner and Thomas, 2010), such as combining direct finger input with deformable physical devices to enable clay-like sculpting operations, including deforming, smoothing, pasting, and extruding (Sheng et al., 2006). Efforts to adapt virtual clay modeling to VR environments have been limited. For instance, Moo-Young et al. (2021) made preliminary attempts to integrate these techniques into VR; however, their approach faced challenges in achieving the robustness and accuracy needed for intricate hand-object interactions. To address this gap, our work introduces a realistic, hands-on modeling framework for VR, leveraging the Material Point Method (Jiang et al., 2016) and efficient hand-object contact handling to enable intuitive and detailed clay-like modeling in virtual environments.
@@
Creative tools for 3D modeling in VR encompass different categories. Among drawing-based modeling tools, Surface Drawing (Schkolne et al., 2001) stands as a pioneering approach, enabling users to draw strokes in space using hand motions and edit 3D models with tangible tools. SurfaceBrush (Rosales et al., 2019) and AdaptiBrush (Rosales et al., 2021) convert dense collections of artist-drawn stroke ribbons into user-intended manifold free-form 3D surfaces. Cassie (Yu et al., 2021) is a conceptual modeling system in VR that leverages free-hand mid-air sketching and a novel 3D optimization framework to create a connected curve network. In addition, inspired by FiberMesh (Nealen et al., 2007), RodMesh (Schulz et al., 2019) replaces the curve drawing with two-handed bending and stretching of virtual rods, allowing users to define outline shapes that are subsequently inflated into manifold mesh surfaces. Commercial tools, such as Shapelab, Adobe Substance 3D Modeler [^1] https://www.adobe.com/products/substance3d/apps/modeler.html, and Kodon [^2] https://www.kodon.xyz/, mainly focus on geometric editing through polygonal or voxel-based sculpting. These tools allow users to manipulate geometry regions and apply operations with dynamic topology updates. Notably, Gravity Sketch introduces subdivision modeling, enabling users to create complex meshes from simple base topology while maintaining smoothness. Inspired by these prior works, we aim to further reduce the demand for professional 3D modeling expertise with physics-aware hands-on shape modeling, making 3D modeling in VR more accessible to novice users.
@@ 4.4. System Implementation
We developed our system using Unity, integrating 3D Gaussian Splatting and Marching Cubes techniques. For the simulation component, we utilized Taichi (Hu et al., 2019) as the compilation engine, which is well-suited for developing high-performance physics-based simulators. The simulation code was compiled into binary files and imported into Unity, where it was executed via C# bindings. For experiments, the system ran on a 16-core 4.3GHz AMD Ryzen 9 9950X machine with an Nvidia RTX 4090 GPU, paired with a Meta Quest 3 VR headset featuring hand-tracking functionality as the display and user input device. This setup enabled efficient and responsive interactions within the VR environment. We will open-source our code and data upon acceptance of this paper.
@@ 5. Evaluation
In this section, we evaluate the real-time performance, visual consistency, and realism of our system, highlighting the contributions of our technical innovations with ablation studies. Unless stated otherwise, the experimental setup is as follows: The simulation uses a grid resolution of $64^{3}$ with 8 particles per grid cell at initialization. Each frame performs 5 simulation substeps, employing a neo-Hookean elasticity model combined with a von Mises plasticity model. For rendering, the density field is sampled at $128^{3}$, with mesh reconstruction via the Marching Cubes algorithm and 5 iterations of Laplacian smoothing to improve surface quality. To ensure consistent comparisons, we recorded a specific hand trajectory for each example and applied it across all tests with varying parameters.
@@
We evaluated the effectiveness of our Uniform Gaussian Volume Regularization by simulating a textured sand cube represented with 3D Gaussians as it was dropped onto the ground under gravity, resulting in significant deformations. Figure 6 compares the original model (left) with the deformed results using (middle) and without (right) the proposed volume regularization loss. Without the loss, the exposed interior reveals noticeable blurry artifacts caused by uneven Gaussian volumes. In contrast, applying the regularization maintains consistent Gaussian volumes, preserving sharp details and achieving high visual fidelity even under extreme deformation.
@@ 6. Case Study
Figure 4 highlights featured modeling operations in VR-Doh, including contact-based shape editing with hands and tools like slabs and rods, mid-air gesture-based bending and twisting, and the sourcing tool for adding new materials. The objects exhibit realistic elastoplastic deformations, closely resembling real-world behavior. This physical realism enables users to intuitively model objects by drawing on their real-life experiences.
@@
Figure 11 demonstrates 8 examples of editing GS-represented objects, showcasing the intuitive and creative capabilities of VR-Doh. First, multiple mushroom-shaped houses were assembled, with their roofs sharpened and the mushrooms enlarged through hand manipulation and mid-air pinch gestures. The text ”VR-Doh” was then sculpted onto the roof by hand. Next, a red pig transitioned from running to performing ballet through flexible selection and dragging of specific body parts, with its head rotated to face the audience using the pinch gesture. Two statues were brought to ’life’: a terracotta figurine joyfully lifted its head and began drumming, while the Discobolus threw its discus with dynamic force. This involved using hands to cut and separate the discus before repositioning it. Moving on, wooden frames were interlocked and freely stacked into an arch through gravity and contact interactions, then merged seamlessly with greenery. The greenery’s complex shapes were also easily refined by hand. A girl’s posture was adjusted as her arms bent to mimic the act of eating a watermelon. Following that, flowers on a vase were flexibly selected using the pinch gesture, then intertwined in space with penetrations automatically avoided, demonstrating precise and convenient control. Finally, a capybara plush toy was made to appear fatter, while its head decoration was reshaped into a strawberry and placed in the front. These examples highlight VR-Doh’s ability to simplify complex modeling tasks while maintaining creativity and controllability. All examples are available on an anonymous online 3D viewer.

### v5

@@ Abstract.
We introduce VR-Doh, an open-source, hands-on 3D modeling system that enables intuitive creation and manipulation of elastoplastic objects in Virtual Reality (VR). By customizing the Material Point Method (MPM) for real-time simulation of hand-induced large deformations and enhancing 3D Gaussian Splatting for seamless rendering, VR-Doh provides an interactive and immersive 3D modeling experience. Users can naturally sculpt, deform, and edit objects through both contact- and gesture-based hand-object interactions. To achieve real-time performance, our system incorporates localized simulation techniques, particle-level collision handling, and the decoupling of physical and appearance representations, ensuring smooth and responsive interactions. VR-Doh supports both object creation and editing, enabling diverse modeling tasks such as designing food items, characters, and interlocking structures, all resulting in simulation-ready assets. User studies with both novice and experienced participants highlight the system’s intuitive design, immersive feedback, and creative potential. Compared to existing geometric modeling tools, VR-Doh offers enhanced accessibility and natural interaction, making it a powerful tool for creative exploration in VR.
@@
submissionid: 318journal: TOGjournalyear: 2025journalvolume: 44journalnumber: 4publicationmonth: 8doi: 10.1145/3731154isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Virtual reality
@@ 1. Introduction
Meanwhile, 3D modeling in Virtual Reality (VR) has become increasingly popular due to its ability to provide immersion and depth, thereby enhancing designers’ creativity and collaboration (Rosales et al., 2019; Yu et al., 2021). Commercial tools like Shapelab [^1] https://shapelabvr.com/ and Gravity Sketch [^2] https://gravitysketch.com/ have advanced this vision. However, their approaches are still limited to procedural methods, such as drawing curves to form surfaces from scratch, which are restricted to geometric techniques and require significant skills.
@@
Building on these foundational advantages, we present VR-Doh, an immersive VR modeling system that integrates physics-based simulation with natural hand interactions to achieve intuitive 3D content creation. Our system addresses two fundamental challenges: 1) enabling physically realistic material deformations through direct hand manipulation, and 2) maintaining real-time performance for responsive interaction despite complex physics computations.
The foundation of VR-Doh lies in our customization of the Material Point Method (MPM) (Jiang et al., 2016; Hu et al., 2018), a hybrid Lagrangian-Eulerian simulation framework with great potential for realistic, interactive modeling. We leverage MPM’s efficacy in handling large elastoplastic deformations to enable natural hand-object interactions, closely mimicking real-world clay manipulation. Furthermore, its particle-based formulation integrates naturally with modern scene representations such as 3D Gaussian Splatting (GS) (Kerbl et al., 2023), allowing VR-Doh to support not only freeform modeling from scratch but also editing of existing photorealistic 3D scenes via extensions of PhysGaussian (Xie et al., 2024). To address the challenges of standard MPM for real-time performance, we introduce three synergistic optimizations: 1) Particle-level collision handling resolves fine geometric details during intensive hand-object interactions without requiring high-resolution grids; 2) Decoupled physical and appearance representations allow detailed 3D GS rendering to be driven by lightweight simulation proxies; and 3) Localized simulation dynamically concentrates computational effort in regions of active manipulation, ensuring responsive performance. Complementing the technical innovations, we design an interaction toolkit that maps hand motions to a wide range of modeling operations. Users can control deformation tools, such as rods and planar slabs, for more structured edits, while mid-air gestures enable precise selection and manipulation via force fields. This multi-modal interaction paradigm broadens creative possibilities while keeping interactions grounded in real-world physics.
To evaluate VR-Doh’s usability and effectiveness, we conducted user studies with participants of varying expertise. The study highlights its intuitive design, immersive feedback, and potential for complementing traditional modeling software in creative exploration. These findings offer insights into how users naturally leverage physics-based interactions, paving the way for future improvements and more immersive 3D modeling experiences in VR.
@@
A key advantage of our approach is its ability to make 3D modeling in VR significantly more intuitive and accessible, especially for novice users. By incorporating physics-based deformations, VR-Doh provides a realistic and immersive modeling experience that surpasses traditional geometric modeling approaches. Hands-on input lowers the barrier to entry, enabling users to interact with virtual objects naturally and intuitively, while flexible selection of operation areas improves efficiency and enhances immersion. This direct interaction mimics real-world manipulation, allowing users to predict outcomes based on everyday experiences and reducing the reliance on specialized skills and technical expertise. VR-Doh is open-sourced at https://github.com/Simulation-Intelligence/VR-Doh.
@@ 2. Related Work
Virtual modeling of clay-like materials using dexterous input offers significant advantages over traditional 3D modeling, including high expressivity and a lower learning curve, effectively replicating the tactile experience of working with physical clay (Sheng et al., 2006; Chatterjee, 2024). Despite these benefits, the computational demands of simulating dexterous hand-object interactions have limited the integration of virtual clay modeling into VR environments, with most existing implementations constrained to 2D screens. Early approaches to virtual clay modeling utilized dynamic subdivision solids (McDonnell et al., 2001) and sculpting tools (Galyean and Hughes, 1991), followed by the application of plasticity models to capture essential physical properties of real clay, such as mass conservation and surface tension effects (Dewaele and Cani, 2004). To improve interactivity, Barreiro et al. (2021) introduced a particle-based viscoplasticity model with ultrasound haptic feedback, although the fidelity of finger-based interactions remained insufficient for detailed 3D modeling tasks. Additionally, physical proxies have been employed to enhance the virtual clay experience (Sheng et al., 2006; Pihuit et al., 2008; Marner and Thomas, 2010), such as combining direct finger input with deformable physical devices to enable clay-like sculpting operations, including deforming, smoothing, pasting, and extruding (Sheng et al., 2006). Efforts to adapt virtual clay modeling to VR environments have been limited. For instance, Jang et al. (2014) and Moo-Young et al. (2021) made preliminary attempts to integrate these techniques into VR; however, their approach faced challenges in achieving the robustness and accuracy needed for intricate hand-object interactions. To overcome these challenges, our work introduces an intuitive, contact-based VR modeling framework leveraging MPM (Jiang et al., 2016) and 3D GS (Kerbl et al., 2023). Our approach significantly enhances real-time performance and hand-object interaction accuracy, enabling a more natural and immersive experience that underscores the unique benefits and potential of clay-like modeling in virtual environments.
@@
Creative tools for 3D modeling in VR encompass different categories. Among drawing-based modeling tools, Surface Drawing (Schkolne et al., 2001) stands as a pioneering approach, enabling users to draw strokes in space using hand motions and edit 3D models with tangible tools. To enable smoother 3D curve creation, Dynamic Dragging (Keefe et al., 2008) introduces an adaptive drag line that adjusts based on curvature and drawing speed. SurfaceBrush (Rosales et al., 2019) and AdaptiBrush (Rosales et al., 2021) convert dense collections of artist-drawn stroke ribbons into user-intended manifold free-form 3D surfaces. Cassie (Yu et al., 2021) is a conceptual modeling system in VR that leverages free-hand mid-air sketching and a novel 3D optimization framework to create a connected curve network. In addition, inspired by FiberMesh (Nealen et al., 2007), RodMesh (Schulz et al., 2019) replaces the curve drawing with two-handed bending and stretching of virtual rods, allowing users to define outline shapes that are subsequently inflated into manifold mesh surfaces. Peng et al. (2020) presents a keyframe-based animated sculpting system that autocompletes user edits through an intuitive brushing interface. Commercial tools, such as Shapelab, Adobe Substance 3D Modeler [^3] https://www.adobe.com/products/substance3d/apps/modeler.html, and Kodon [^4] https://www.kodon.xyz/, mainly focus on geometric editing through polygonal or voxel-based sculpting. These tools allow users to manipulate geometry regions and apply operations with dynamic topology updates. Notably, Gravity Sketch introduces subdivision modeling, enabling users to create complex meshes from simple base topology while maintaining smoothness. Inspired by these prior works, we aim to further reduce the demand for professional 3D modeling expertise with physics-aware hands-on shape modeling, making 3D modeling in VR more accessible to novice users.
@@ 4.1. Hand-based User Interaction
Figure 3. Medial axis approximation of hand geometry. The highlighted curves indicate the boundaries of a specific medial slab and a medial cone.
@@
Figure 4. Featured Operations: (a) Pinching and reshaping with fingers; (b) Using a sourcing tool to extrude new materials onto an object; (c) Mid-air gestures for bending and twisting objects; (d) Hand-based cutting of objects; (e) (f) Tool-based operations with a slab and a rod; .
@@ 4.4. System Implementation
We developed our system using Unity, integrating 3D Gaussian Splatting and Marching Cubes techniques. For the simulation component, we utilized Taichi (Hu et al., 2019) as the compilation engine, which is well-suited for developing high-performance physics-based simulators. The simulation code was compiled into binary files and imported into Unity, where it was executed via C# bindings. For experiments, the system ran on a 16-core 4.3GHz AMD Ryzen 9 9950X machine with an Nvidia RTX 4090 GPU, paired with a Meta Quest 3 VR headset featuring hand-tracking functionality as the display and user input device. This setup enabled efficient and responsive interactions within the VR environment.
@@ 5. Evaluation
In this section, we evaluate the real-time performance, visual consistency, and realism of our system, highlighting the contributions of our technical innovations with ablation studies. Unless stated otherwise, the experimental setup is as follows: The simulation uses a grid resolution of $64^{3}$ with 8 particles per grid cell at initialization. Each frame performs 5 simulation substeps, employing a neo-Hookean elasticity model (Young’s modulus $E=10^{6}Pa$, Poisson’s ratio $\nu=0.3$) combined with a von Mises plasticity model (yield stress $\sigma_{y}=10^{5}Pa$). For rendering, the density field is sampled at $128^{3}$, with mesh reconstruction via the Marching Cubes algorithm and 5 iterations of Laplacian smoothing to improve surface quality. To ensure consistent comparisons, we recorded a specific hand trajectory for each example and applied it across all tests with varying parameters.
@@
Figure 5. Real-Time Performance Evaluation. FPS of VR-Doh with varying simulation grid resolution, substeps per frame, and MPM particles per cell in a sphere compression scenario (Figure 7, bottom). Results demonstrate VR-Doh’s ability to maintain real-time performance under high computational demands.
@@
Figure 6. Convergence under Spatial Refinement. Sphere deformations under identical hand imprints and squeezing converge at resolutions above $48^{3}$, ensuring accuracy without excessive computational cost.
Figure 7. Particle-Level Collision Handling. Comparison of MPM boundary conditions with and without particle-level collision handling on a hand imprint (top) and a sphere compression (bottom) example. Combining both methods (right) achieves realistic deformations with enhanced detail while avoiding volume loss.
@@
We evaluated the effectiveness of our Uniform Gaussian Volume Regularization by simulating a textured sand cube represented with 3D Gaussians as it was dropped onto the ground under gravity, resulting in significant deformations. Figure 8 compares the original model (left) with the deformed results using (middle) and without (right) the proposed volume regularization loss, both trained for 30,000 iterations using Adam (Kingma and Ba, 2014). Without the loss, the exposed interior reveals noticeable blurry artifacts caused by uneven Gaussian volumes. In contrast, applying the regularization maintains consistent Gaussian volumes, preserving sharp details and achieving high visual fidelity even under extreme deformation.
Figure 8. Uniform Gaussian Volume Regularizer. Comparison of a sand cube dropped under gravity shows that applying the regularization (middle) preserves sharp details and avoids blurry artifacts in the interior (right), ensuring high visual fidelity during large deformations.
@@
Figure 9. Decoupled Appearance and Physical Representations. This example of a wolf sitting under gravity demonstrates that reducing the number of MPM particles driving the 3D Gaussians slightly increases the wolf’s volume but maintains high visual fidelity while significantly improving FPS, showcasing the effectiveness of our approach.
@@ 6. Case Study
Figure 4 highlights featured modeling operations in VR-Doh, including contact-based shape editing with hands and tools like slabs and rods, mid-air gesture-based bending and twisting, and the sourcing tool for adding new materials. The objects exhibit realistic elastoplastic deformations, closely resembling real-world behavior. This physical realism enables users to intuitively model objects by drawing on their real-life experiences. Users can freely adjust material parameters such as stiffness and yield stress, or choose from a set of predefined materials to suit different modeling needs.
@@
Figure 10. Object Creation Examples. Crafting a snowman, assembling a hamburger with gravity-stacked layers, creating a panda holding bamboo, rolling a Swiss roll, and shaping steamed buns in a bamboo steamer, showcasing intuitive and precise 3D modeling.
@@
Figure 11 demonstrates 8 examples of editing GS-represented objects, showcasing the intuitive and creative capabilities of VR-Doh. First, multiple mushroom-shaped houses were assembled, with their roofs sharpened and the mushrooms enlarged through hand manipulation and mid-air pinch gestures. The text ”VR-Doh” was then sculpted onto the roof by hand. Next, a red pig transitioned from running to performing ballet through flexible selection and dragging of specific body parts, with its head rotated to face the audience using the pinch gesture. Two statues were brought to ”life”: a terracotta figurine joyfully lifted its head and began drumming, while the Discobolus threw its discus with dynamic force. This involved using hands to cut and separate the discus before repositioning it. Subsequently, wooden frames were interlocked and freely stacked into an arch through gravity and contact interactions, then merged seamlessly with greenery. The greenery’s complex shapes were also easily refined by hand. A girl’s posture was adjusted as her arms bent to mimic the act of eating a watermelon. Following that, flowers on a vase were flexibly selected using the pinch gesture, then intertwined in space with penetrations automatically avoided, demonstrating precise and convenient control. Finally, a capybara plush toy was made to appear fatter, while its head decoration was reshaped into a strawberry and placed in the front. These examples highlight VR-Doh’s ability to simplify complex modeling tasks while maintaining creativity and controllability.
Figure 11. 3D GS object editing examples. Users combined, reshaped, and customized models using hand gestures and tools, bringing objects to life, merging elements seamlessly, and refining intricate details with ease. These examples highlight VR-Doh’s ability to simplify complex modeling tasks while preserving creativity and control.